OpenStudy (anonymous):

(adj B)^-1 = adj(B^-1)

OpenStudy (anonymous):

( adj B)^-1 =\[\frac{ 1 }{ \det B } B\]

OpenStudy (experimentx):

??

OpenStudy (zarkon):

are you allowed to use \[A^{-1}=\frac{1}{\det(A)}\text{adj}(A)\]

OpenStudy (anonymous):

yess

OpenStudy (experimentx):

sorry OS went down ... use the above defn ... you will get \( |B| B = |B|B\)

OpenStudy (zarkon):

\[\text{adj}(A)=\det(A)A^{-1}\] \[(\text{adj}(A))^{-1}=(\det(A)A^{-1})^{-1}=(\det(A))^{-1}(A^{-1})^{-1}=\frac{1}{\det(A)}A\]

OpenStudy (anonymous):

how about this?

OpenStudy (zarkon):

Replace the A's above by B's then use \[\text{adj}(B)=\det(B)B^{-1}\] replace \(B\) with \(B^{-1}\)

OpenStudy (anonymous):

imean by the other one....

OpenStudy (zarkon):

what do you mean by "by the other one..."?

OpenStudy (anonymous):

here

OpenStudy (anonymous):

see the attachment please

OpenStudy (zarkon):

yes...I told you what you could do to solve that

OpenStudy (anonymous):

??

OpenStudy (zarkon):

use \[\text{adj}(B)=\det(B)B^{-1}\] replace \(B\) by \(B^{-1}\). What do you get

OpenStudy (anonymous):

dont know

OpenStudy (zarkon):

\[\text{adj}(B^{-1})=\det(B^{-1})(B^{-1})^{-1}\] simplify this (the right hand side)

OpenStudy (anonymous):

I??

OpenStudy (zarkon):

no

OpenStudy (anonymous):

then?

OpenStudy (zarkon):

if you can't finish from here then you need to go and review the properties of invertible matrices

OpenStudy (zarkon):

I've practically done the entire problem

OpenStudy (anonymous):

what i get if adj.adj A

OpenStudy (turingtest):

I don't even see what that means. @Zarkon 's proof is almost complete and totally clear. If you don't understand any of it you really do need to review matrix properties as he said.